OpenStudy (anonymous):

a snowflake weighs 0.000001 gram what is the weight as a power of 10.

OpenStudy (anonymous):

0.000001 = 1 / 1000000 = \[\frac{ 1 }{ 10^{6} } = 1 \times 10^{-6}\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

you are most welcomed @kathyforever